OpenStudy (anonymous):

Can someone please answer this question. Solve for x. Log(x+2)+log(x-4)=log55

OpenStudy (therealmeeeee):

2 log 55/26 + log 28/175 =2(log 55 − log 26) +log28 −log 175 2log(11×5)− 2log (2×13) +(4×7)−log (25×7) =2(log11 +log5)− 2(log 2 +log 13) + log4 +log7−(log 25+log7) =2log11 + 2log5− 2log 2 −2log 13 + log4 +log7−log 25−log7 =2log11 −2log 13 + 2log 5− 2log 2 + log2² +log7−log 5²−log7 =2log11 −2log 13 + 2log 5− 2log 2 +2log2 +log7−2log 5−log7 =2log11 −2log 13 + 2log 5 −2log 5− 2log 2 +2log2 +log7−log7 =2log11 −2log 13 =2x−2y =2(x−y)

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

No because x is not isolated

OpenStudy (therealmeeeee):

ohhh